OpenStudy (anonymous):

Solve: 3(x + 4) = 5(x − 2) 11 −11 1 −1

OpenStudy (theviper):

3(x+4)=5(x-2) 3x+12=5x-10 5x-3x=12+10 2x=22 x=22/2=11 So 11 is the right answer ;)

OpenStudy (theviper):

Got it?

OpenStudy (anonymous):

Yes Thank Your Explaining